# On the Effectiveness and Generalization of Race Representations for Debiasing High-Stakes Decisions

### v1

@@ 7 Conclusion
In this work, we found that prompting alone is insufficient to reduce models’ biases, and in many cases can worsen them or completely derail model behavior. In contrast, representation-based debiasing is a more promising approach, as we found subspaces of model hidden representations that strongly encode race and intervened on them to reduce Gemma’s bias by 37-57%. Our work also surprisingly discovered that race representations may not generalize to different task settings, as a seemingly innocuous change in the prompt template can result in a different race representation. We believe more research is needed to understand how and why LLMs’ race representations differ across task settings.
@@
Our work focuses on understanding and mitigating biases of LLMs in high-stakes decision making, which contributes towards the goal of responsible AI. However, it is useful to point out that although the race subspace can be used for debiasing, it could also be used to manipulate model decisions in an undesirable fashion. In fact, our results show that prompting the model to be fair can already be used as a way to achieve extreme disparities, highlighting a way to negatively steer models undetectable by human intuition.
@@ Appendix A Datasets and prompts
We show examples of a few notable prompt formats. We mostly show Gemma prompts to illustrate the main idea since LLaMA prompts are mostly identical to Gemma prompts except for a few minor details such as changing the specified acceptance rate in order to prevent the model from accepting or rejecting too many applicants. In general, the prompt should and likely will have to be customized for the model and task combination, since different models might respond differently on the same task, and the same model might respond differently on different tasks. A few challenges we encountered with prompting models include the model accepting everyone, rejecting everyone, or demonstrates no bias. It is surprising how sensitive LLMs can be to the prompt format.
@@ Appendix C Alignment training
Target Source Original label Counter-factual label
… should a college admissions officer admit Scott to Harvard University? The applicant has a 3.46 out of 4.0 GPA, participates in 8 impactful on-campus clubs, and has 2 strong recommendation letters… … should a college admissions officer admit Deandre to Northwestern University? The applicant has a 1.61 out of 4.0 GPA, participates in 0 impactful on-campus clubs, and has 1 strong recommendation letters… Yes No
.. should a college admissions officer admit Gustavo to University of Michigan? The applicant has a 3.69 out of 4.0 GPA, participates in 7 impactful on-campus clubs, and has 3 strong recommendation letters… … should a college admissions officer admit Mei to Harvard University? The applicant has a 1.73 out of 4.0 GPA, participates in 2 impactful on-campus clubs, and has 0 strong recommendation letters… No Yes
Table 10: Examples of interchange interventions from the Admissions counterfactual dataset.
@@ Appendix E Race subspace generalization
Figure 15: There is a lack of generalization between race subspaces trained on different decision settings. $S\to T$ denotes an intervention from a source domain $S$ to a target domain $T$. The results are averaged over layers 10 and 11.

### v2

@@ 7 Conclusion
In this work, we found that prompting alone is insufficient to reduce models’ biases, and in many cases can worsen them or completely derail model behavior. In contrast, representation-based debiasing is a more promising approach, as we found subspaces of models’ hidden representations that strongly encode race and intervened on them to reduce Gemma’s bias by 37-57%. Our work also surprisingly discovered that race representations may not generalize to different task settings, as a seemingly innocuous change in the prompt template can result in a different race representation. We believe more research is needed to understand how and why LLMs’ race representations differ across task settings.
@@
Using name as a proxy for race.
In the NLP literature on inferring demographic attributes from name, several authors have found that this approach may have limited construct validity (Gautam et al., 2024). For certain ethnicities, names may have strong race associations, while for others, the associations may be weaker. We believe this issue is mitigated by An et al. (2024)’s data collection procedure: 1) For each name, classify its associated race and gender by selecting the majority (>>50%) classes. 2) For each race-gender subgroup, e.g., Black female, they select the top 50 ranked by the percentage that race.
Nevertheless, this majority-class assignment approach is still prone to misclassification, as noted by Gautam et al. (2024). However, since our paper is not primarily concerned with predicting people’s races based on names, but with removing any existing association, we believe that there is limited harm. Our datasets can serve as a starting point for benchmarking debiasing methods.
Dual use concern.
Our work focuses on understanding and mitigating biases of LLMs in high-stakes decision making, which contributes towards the goal of responsible AI. However, it is noteworthy that although the race subspace can be used for debiasing, it could also be used to manipulate model decisions in an undesirable fashion, an ethical concern known as dual use. This concern may be mitigated with proprietary models, where an attacker might not have access to the model’s activations, but our results with prompting also show that prompting the model for fairness can be used as a way to achieve extreme disparities. Although this is not inherent to our approach, it highlights that in addition to the well-recognized concern that large language models can be negatively steered, they may lead to unintentional adverse effects even when used with good intentions.
@@ Appendix A Datasets and prompts
We show examples of a few notable prompt formats. We mostly show Gemma prompts to illustrate the main idea since LLaMA prompts are mostly identical to Gemma prompts except for a few minor details such as changing the specified acceptance rate in order to prevent the model from accepting or rejecting too many applicants. In general, the prompt should and likely will have to be customized for the model and task combination, since different models might respond differently on the same task, and the same model might respond differently on different tasks. A few challenges we encountered with prompting models include the model accepting everyone, rejecting everyone, or demonstrating no bias. It is surprising how sensitive LLMs can be to the prompt format.
@@ Appendix B Models’ biases
Method Bias Score $\downarrow$ Outcome $\Delta$ (%) Asian Black Latino White
Original 13.54 0.00 58.00 49.75 51.75 62.00
No Name 0.00 9.38 64.75 64.75 64.75 64.75
Prompting Fewshot 5.11 35.38 91.50 88.25 90.25 93.00
Prompting CoT 21.65 -35.00 19.75 18.75 19.75 23.25
Table 11: Few-shot and chain-of-thought prompting fails to correctly debias Gemma in Admissions.
Due to a lack of groundtruth decisions, i.e., for a given applicant profile, we have no prior over their likelihood of acceptance, we cannot properly include few-shot examples in the prompt. To work around this, we append 8 examples to all prompts in Admissions, 4 cases of acceptance, and 4 of rejection. In each group, each sample differs only by their race, and their decision are shown to be the same.
While few-shot prompting reduces the bias, it drastically increases the average acceptance rate (Table 11). In contrast, chain-of-thought prompting exacerbates the bias while significantly decreasing the acceptance rate. Hence, despite their success in increasing performance in the AI literature, neither method are effective at controlling models when they are strongly biased.
@@ Appendix E Race subspace generalization
Figure 15: There is a lack of generalization between race subspaces trained on different decision settings. $S\to T$ denotes an intervention from a source domain $S$ to a target domain $T$. The results are averaged over layers 10 and $1$.
@@
Layer Original Race Avg Race Proj
11 0.88 0.50 0.44
12 0.87 0.51 0.45
13 0.80 0.50 0.46
14 0.79 0.48 0.45
15 0.72 0.49 0.45
16 0.80 0.50 0.46
17 0.87 0.00 0.00
Table 16: Searching for a race subspace after performing debiasing interventions at layer 10.